# Why Machine Learning Models Systematically Underestimate Extreme Values

### v1

@@ Appendix B Variance of Linear Coefficient Estimator
$\hat{\beta}=\frac{\sum_{i}x_{\text{obs},i}y_{\text{obs},i}}{\sum_{i}x_{\text{obs},i}^{2}}$ (B1)
@@
$\hat{\beta}=\frac{\sum_{i}x_{\text{obs},i}(\beta x_{\text{obs},i}-\beta\delta_{x,i}+\delta_{y,i})}{\sum_{i}x_{\text{obs},i}^{2}}$ (B2)
@@
$\hat{\beta}=\beta+\frac{\sum_{i}x_{\text{obs},i}(-\beta\delta_{x,i}+\delta_{y,i})}{\sum_{i}x_{\text{obs},i}^{2}}$ (B3)
@@
$\operatorname{Var}(\hat{\beta})=\operatorname{Var}\left(\frac{\sum_{i}x_{\text{obs},i}(-\beta\delta_{x,i}+\delta_{y,i})}{\sum_{i}x_{\text{obs},i}^{2}}\right)$ (B4)
@@
$=\frac{\sum_{i}x_{\text{obs},i}^{2}}{(\sum_{i}x_{\text{obs},i}^{2})^{2}}(\sigma_{y}^{2}+\beta^{2}\sigma_{x}^{2})$ (B5)
@@
$\operatorname{Var}(\hat{\beta})=\frac{\sigma_{y}^{2}+\beta^{2}\sigma_{x}^{2}}{n(\sigma_{\text{range}}^{2}+\sigma_{x}^{2})}.$ (B6)
@@
$\hat{\beta}_{j}=\frac{\sum_{i}x_{\text{obs},ij}(\sum_{k=1}^{p}\beta_{k}x_{\text{true},ik}+\delta_{y,i})}{\sum_{i}x_{\text{obs},ij}^{2}}$ (B7)
@@
$=\beta_{j}+\frac{\sum_{i}x_{\text{obs},ij}(-\beta_{j}\delta_{x,ij}+\delta_{y,i}+\sum_{k\neq j}\beta_{k}x_{\text{true},ik})}{\sum_{i}x_{\text{obs},ij}^{2}}$ (B8)
@@
$\operatorname{Var}(\hat{\beta}_{j})$ $=\sum_{i}\left(\frac{x_{\text{obs},ij}}{\sum_{\ell}x_{\text{obs},\ell j}^{2}}\right)^{2}\operatorname{Var}(-\beta_{j}\delta_{x,ij}+\delta_{y,i}+\sum_{k\neq j}\beta_{k}x_{\text{true},ik})$ (B9)
@@
$\operatorname{Var}(-\beta_{j}\delta_{x,ij}+\delta_{y,i}+\sum_{k\neq j}\beta_{k}x_{\text{true},ik})$ $=\beta_{j}^{2}\sigma_{x}^{2}+\sigma_{y}^{2}+\sum_{k\neq j}\beta_{k}^{2}\operatorname{Var}(x_{\text{true},ik})$ (B10)
The crucial step comes in evaluating $\operatorname{Var}(x_{\text{true},ik})$ conditional on $x_{\text{obs}}$:
@@
$\operatorname{Var}(\hat{\beta}_{j})$ $=\frac{\sum_{i}x_{\text{obs},ij}^{2}}{(\sum_{i}x_{\text{obs},ij}^{2})^{2}}(\sigma_{y}^{2}+\sum_{k=1}^{p}\beta_{k}^{2}\sigma_{x}^{2})$ (B12)
@@
$\operatorname{Var}(\hat{\beta}_{j})=\frac{\sigma_{y}^{2}+\sum_{k=1}^{p}\beta_{k}^{2}\sigma_{x}^{2}}{n(\sigma_{\text{range}}^{2}+\sigma_{x}^{2})}$ (B13)
@@ Appendix C Derivation of the attenuation bias for correlated multivariate features
$\mathbb{E}[\hat{\boldsymbol{\beta}}]\approx(\mathbb{E}[\mathbf{X}_{\text{true}}^{\top}\mathbf{X}_{\text{true}}]+n\sigma_{x}^{2}\mathbf{I}_{p})^{-1}\mathbb{E}[\mathbf{X}_{\text{true}}^{\top}\mathbf{X}_{\text{true}}]\boldsymbol{\beta}$ (C1)
@@
$\mathbf{X}_{\text{true}}=x_{\text{true}}\mathbf{A},$ (C2)
@@
$=n\,\mathbb{E}[x_{\text{true}}^{2}]\mathbf{A}^{\top}\mathbf{A}.$ (C3)
@@
$\mathbb{E}[\mathbf{X}_{\text{true}}^{\top}\mathbf{X}_{\text{true}}]\approx S_{x}\mathbf{A}^{\top}\mathbf{A}.$ (C4)
Substituting these expressions into Equation (C1), we have:
@@
$=\left(\mathbf{A}^{\top}\mathbf{A}+\frac{n\sigma_{x}^{2}}{S_{x}}\mathbf{I}_{p}\right)^{-1}\mathbf{A}^{\top}\mathbf{A}\boldsymbol{\beta}.$ (C5)
@@
$\mathbf{A}^{\top}\mathbf{A}=\mathbf{a}\mathbf{a}^{\top}.$ (C6)
@@
$\mathbf{M}=\mathbf{a}\mathbf{a}^{\top}+\kappa\mathbf{I}_{p}.$ (C7)
@@
$(\mathbf{a}\mathbf{a}^{\top}+\kappa\mathbf{I}_{p})^{-1}=\kappa^{-1}\left(\mathbf{I}_{p}-\frac{\mathbf{a}\mathbf{a}^{\top}}{\kappa+\mathbf{a}^{\top}\mathbf{a}}\right).$ (C8)
@@
$=\kappa^{-1}\left(\mathbf{a}\mathbf{a}^{\top}\boldsymbol{\beta}-\frac{\mathbf{a}\mathbf{a}^{\top}\mathbf{a}\mathbf{a}^{\top}}{\kappa+\mathbf{a}^{\top}\mathbf{a}}\boldsymbol{\beta}\right).$ (C9)
@@
$\mathbf{a}\mathbf{a}^{\top}\mathbf{a}=(\mathbf{a}^{\top}\mathbf{a})\mathbf{a}.$ (C10)
@@
$=\kappa^{-1}(\mathbf{a}^{\top}\boldsymbol{\beta})\mathbf{a}\left(1-\frac{\mathbf{a}^{\top}\mathbf{a}}{\kappa+\mathbf{a}^{\top}\mathbf{a}}\right).$ (C11)
@@
$1-\frac{\mathbf{a}^{\top}\mathbf{a}}{\kappa+\mathbf{a}^{\top}\mathbf{a}}$ $=\frac{\kappa}{\kappa+\mathbf{a}^{\top}\mathbf{a}}.$ (C12)
@@
$=\frac{\mathbf{a}^{\top}\boldsymbol{\beta}}{\kappa+\mathbf{a}^{\top}\mathbf{a}}\mathbf{a}$ (C13)
$\equiv\lambda_{p}\mathbf{a}.$ (C14)
@@
$\lambda_{p}=\frac{\mathbf{a}^{\top}\boldsymbol{\beta}}{\kappa+\mathbf{a}^{\top}\mathbf{a}}.$ (C15)
@@
$\mathbb{E}[\hat{\boldsymbol{\beta}}]\approx\frac{\beta\mathbf{a}^{\top}\mathbf{a}}{\kappa+\mathbf{a}^{\top}\mathbf{a}}\mathbf{a}.$ (C16)
@@
$\mathbb{E}[\hat{\beta}_{j}]\approx\lambda_{p}\beta a_{j}=\lambda_{p}\beta_{j},$ (C17)
and we arrive at the analytic expression for the attenuation factor $\lambda_{p}$:
$\lambda_{p}=\frac{\mathbf{a}^{\top}\mathbf{a}}{\kappa+\mathbf{a}^{\top}\mathbf{a}}=\frac{\sum_{j=1}^{p}a_{j}^{2}}{\frac{n\sigma_{x}^{2}}{S_{x}}+\sum_{j=1}^{p}a_{j}^{2}}.$ (C18)

### v2

@@ Appendix B Variance of Linear Coefficient Estimator
$\hat{\beta}=\frac{\sum_{i}x_{\text{obs},i}y_{\text{obs},i}}{\sum_{i}x_{\text{obs},i}^{2}}$ (53)
@@
$\hat{\beta}=\frac{\sum_{i}x_{\text{obs},i}(\beta x_{\text{obs},i}-\beta\delta_{x,i}+\delta_{y,i})}{\sum_{i}x_{\text{obs},i}^{2}}$ (54)
@@
$\hat{\beta}=\beta+\frac{\sum_{i}x_{\text{obs},i}(-\beta\delta_{x,i}+\delta_{y,i})}{\sum_{i}x_{\text{obs},i}^{2}}$ (55)
@@
$\operatorname{Var}(\hat{\beta})=\operatorname{Var}\left(\frac{\sum_{i}x_{\text{obs},i}(-\beta\delta_{x,i}+\delta_{y,i})}{\sum_{i}x_{\text{obs},i}^{2}}\right)$ (56)
@@
$=\frac{\sum_{i}x_{\text{obs},i}^{2}}{(\sum_{i}x_{\text{obs},i}^{2})^{2}}(\sigma_{y}^{2}+\beta^{2}\sigma_{x}^{2})$ (57)
@@
$\hat{\beta}_{j}=\frac{\sum_{i}x_{\text{obs},ij}(\sum_{k=1}^{p}\beta_{k}x_{\text{true},ik}+\delta_{y,i})}{\sum_{i}x_{\text{obs},ij}^{2}}$ (59)
@@
$=\beta_{j}+\frac{\sum_{i}x_{\text{obs},ij}(-\beta_{j}\delta_{x,ij}+\delta_{y,i}+\sum_{k\neq j}\beta_{k}x_{\text{true},ik})}{\sum_{i}x_{\text{obs},ij}^{2}}$ (60)
@@
$\operatorname{Var}(\hat{\beta}_{j})$ $=\sum_{i}\left(\frac{x_{\text{obs},ij}}{\sum_{\ell}x_{\text{obs},\ell j}^{2}}\right)^{2}\operatorname{Var}(-\beta_{j}\delta_{x,ij}+\delta_{y,i}+\sum_{k\neq j}\beta_{k}x_{\text{true},ik})$ (61)
@@
$\operatorname{Var}(-\beta_{j}\delta_{x,ij}+\delta_{y,i}+\sum_{k\neq j}\beta_{k}x_{\text{true},ik})$ $=\beta_{j}^{2}\sigma_{x}^{2}+\sigma_{y}^{2}+\sum_{k\neq j}\beta_{k}^{2}\operatorname{Var}(x_{\text{true},ik})$ (62)
The key step comes in evaluating $\operatorname{Var}(x_{\text{true},ik})$ conditional on $x_{\text{obs}}$:
@@
$\operatorname{Var}(\hat{\beta}_{j})$ $=\frac{\sum_{i}x_{\text{obs},ij}^{2}}{(\sum_{i}x_{\text{obs},ij}^{2})^{2}}(\sigma_{y}^{2}+\sum_{k=1}^{p}\beta_{k}^{2}\sigma_{x}^{2})$ (64)
@@
$\operatorname{Var}(\hat{\beta}_{j})=\frac{\sigma_{y}^{2}+\sum_{k=1}^{p}\beta_{k}^{2}\sigma_{x}^{2}}{n(\sigma_{\text{range}}^{2}+\sigma_{x}^{2})}$ (65)
@@ Appendix C Derivation of the attenuation bias for correlated multivariate features
$\mathbb{E}[\hat{\boldsymbol{\beta}}]\approx(\mathbb{E}[\mathbf{X}_{\text{true}}^{\top}\mathbf{X}_{\text{true}}]+n\sigma_{x}^{2}\mathbf{I}_{p})^{-1}\mathbb{E}[\mathbf{X}_{\text{true}}^{\top}\mathbf{X}_{\text{true}}]\boldsymbol{\beta}$ (66)
@@
$\mathbf{X}_{\text{true}}=x_{\text{true}}\mathbf{A},$ (67)
@@
$=n\,\mathbb{E}[x_{\text{true}}^{2}]\mathbf{A}^{\top}\mathbf{A}.$ (68)
@@
$\mathbb{E}[\mathbf{X}_{\text{true}}^{\top}\mathbf{X}_{\text{true}}]\approx S_{x}\mathbf{A}^{\top}\mathbf{A}.$ (69)
Substituting these expressions into Equation (66), we have:
@@
$=\left(\mathbf{A}^{\top}\mathbf{A}+\frac{n\sigma_{x}^{2}}{S_{x}}\mathbf{I}_{p}\right)^{-1}\mathbf{A}^{\top}\mathbf{A}\boldsymbol{\beta}.$ (70)
@@
$\mathbf{A}^{\top}\mathbf{A}=\mathbf{a}\mathbf{a}^{\top}.$ (71)
@@
$\mathbf{M}=\mathbf{a}\mathbf{a}^{\top}+\kappa\mathbf{I}_{p}.$ (72)
@@
$(\mathbf{a}\mathbf{a}^{\top}+\kappa\mathbf{I}_{p})^{-1}=\kappa^{-1}\left(\mathbf{I}_{p}-\frac{\mathbf{a}\mathbf{a}^{\top}}{\kappa+\mathbf{a}^{\top}\mathbf{a}}\right).$ (73)
@@
$=\kappa^{-1}\left(\mathbf{a}\mathbf{a}^{\top}\boldsymbol{\beta}-\frac{\mathbf{a}\mathbf{a}^{\top}\mathbf{a}\mathbf{a}^{\top}}{\kappa+\mathbf{a}^{\top}\mathbf{a}}\boldsymbol{\beta}\right).$ (74)
@@
$\mathbf{a}\mathbf{a}^{\top}\mathbf{a}=(\mathbf{a}^{\top}\mathbf{a})\mathbf{a}.$ (75)
@@
$=\kappa^{-1}(\mathbf{a}^{\top}\boldsymbol{\beta})\mathbf{a}\left(1-\frac{\mathbf{a}^{\top}\mathbf{a}}{\kappa+\mathbf{a}^{\top}\mathbf{a}}\right).$ (76)
@@
$1-\frac{\mathbf{a}^{\top}\mathbf{a}}{\kappa+\mathbf{a}^{\top}\mathbf{a}}$ $=\frac{\kappa}{\kappa+\mathbf{a}^{\top}\mathbf{a}}.$ (77)
@@
$=\frac{\mathbf{a}^{\top}\boldsymbol{\beta}}{\kappa+\mathbf{a}^{\top}\mathbf{a}}\mathbf{a}$ (78)
$\equiv\lambda_{\beta}\mathbf{a}.$ (79)
@@
$\lambda_{\beta}=\frac{\mathbf{a}^{\top}\boldsymbol{\beta}}{\kappa+\mathbf{a}^{\top}\mathbf{a}}.$ (80)
@@
$\mathbb{E}[\hat{\boldsymbol{\beta}}]\approx\frac{\beta\mathbf{a}^{\top}\mathbf{a}}{\kappa+\mathbf{a}^{\top}\mathbf{a}}\mathbf{a}.$ (81)
@@
$\mathbb{E}[\hat{\beta}_{j}]\approx\lambda_{\beta}\beta a_{j}=\lambda_{\beta}\beta_{j},$ (82)
and we arrive at the analytic expression for the attenuation factor $\lambda_{\beta}$:
$\lambda_{\beta}=\frac{\mathbf{a}^{\top}\mathbf{a}}{\kappa+\mathbf{a}^{\top}\mathbf{a}}=\frac{\sum_{j=1}^{p}a_{j}^{2}}{\frac{n\sigma_{x}^{2}}{S_{x}}+\sum_{j=1}^{p}a_{j}^{2}}.$ (83)